Let X be the time in minutes between two successive arrivals at the drive-up window of a local bank. If X has an exponential distribution with λ-05, compute the following: (If necessary, round your answer to three decimal places.) (a) The expected time between two successive arrivals is (b) The standard deviation of the time between successive arrivals is minutes. minutes. (c) PXs4) (d) P1sXs 3) (e) PX2 1.S D Show My Work (optional

Let X be the time in minutes between two successive arrivals at the drive-up window of a local bank. If X has an exponential distribution with λ=0.5, compute the following: (If necessary, round your answer to three decimal places.) 

(a) The expected time between two successive arrivals is 

(b) The standard deviation of the time between successive arrivals is minutes. minutes. 

(c) P(X≤4) 

(d) P(1≤X≤3) 

(e) P(X≥1.5) 
Answer #1

Here we have exponential distribution with  lambda = 0.5

a. Expected time :

Mear -__ - 0,5

b. Standard deviation :

Variance- = 4 052

Standard deviation = variance V4-2

CDF of exponential distribution is

F(x) =p ( Xleq x ) =   AT

c.

p ( x leq 4 )

= 1-e-0.5*4

= 1 - e-2

= 1 - 0.1353

= 0.8647

d. p ( 1 leq x leq 3 )

= p ( x leq 3 ) - p ( x leq 1 )

= 1-e^{-0.5*3} - 0.5*1

= 0.5*1 053+e

= - 0.2231 + 0.6065

= 0.3834

e) p ( x geq 1.5 )

= 1 - p ( x leq 1.5 )

= 1 - (1-e^{-0.5*1.5})

= e^{-0.5*1.5}

= 0.4724